If a race car can go around a track in 5 minutes,then how long will it take to go around the track if it increases its speed by 20%?
Found 2 solutions by RAY100, Theo:
Answer by RAY100(1637)

v=d/t
.
(1) v=d/5
.
(2) 1.2 v = d/t
.
from (2),,v=d/(1.2 t) = d /5,,,,,,from(1)
.
1.2 t =5
.
t= 5/1.2 = 4.17 min
.
checking,,,,let d=10, v= 10/5 =2
v2= 10/4.17 = 2.4
.
2.4 / 2 = 1.2,,,or 20 % higher,,,ok

Answer by Theo(13342)

the answer should be 5/1.2 = 4.16666666 minutes = 4 minutes and 10 seconds.
here's how I got it.
basic formula is rate * time = distance.
D is going to be the same so we call it D.
The rate will be different so we'll call it R1 and R2.
the time will be different so we'll call it T1 and T2.
we know the following:
R1 * T1 = D
T1 = 5, so
R1 * 5 = D
we know that R2 will be 20% more than R1 so we know that R2 = 1.2 * R1.
we can make a general equation for R2 that says
R2 * T2 = D
R2 will be different.
T2 will be different.
D will remain the same.
since we know that R2 = 1.2 * R1, we can substitute to get
1.2 * R1 * T2 = D
since they both = D, we can make them equal to each other, so we have
R1 * T1 = 1.2 * R1 * T2
since we know that T1 = 5, we can substitute in the equation to get
R1 * 5 = 1.2 * R1 * T2
if we divide both sides of the equation by 1.2 * R1, then we get
5/1.2 = T2
simplifying gets 4.16666666 = T2
taking the fraction and converting it to minutes we get
.1666666666 * 60 = 10
which gets us
4 minutes and 10 seconds.
to check this out, we make an assumption by saying the distance is an arbitrary number such as 10 miles.
In order for this to be done in 5 minutes, the car has to go 10 / 5 minutes = 2 miles a minute which equates to 120 miles an hour.
if we raise 2 miles a minute by 20% we get 2.4 miles a minute which equates to 144 miles an hour.
at 2.4 miles a minute, it takes 10 / 2.4 minutes to complete the distance which equals 4.166666666 minutes which gets us to our answer so we're good.